Shoney Video Concepts produces a line of videodisc players to be linked to personal computers for video games. Videodiscs have much faster access time than tape. With such a computer/video link, the game becomes a very realistic experience. In a simple driving game where the joystick steers the vehicle, for example, rather than seeing computer graphics on the screen, the player is actually viewing a segment of a videodisc shot from a real moving vehicle. Depending on the action of the player (hitting a guard rail, for example), the disc moves virtually instantaneously to that segment and the player becomes part of an actual accident of real vehicles (staged, of course). Shoney is trying to determine a production plan for the next 12 months. The main criterion for this plan is that the employment level is to be held constant over the period. Shoney is continuing in its R&D efforts to develop new applications and prefers not to cause any adverse feelings with the local workforce. For the same reason, all employees should put in full workweeks, even if this is not the lowest-cost alternative. The forecast for the next 12 months is

MONTH FORECAST DEMAND MONTH FORECAST DEMAND
January 550 July 150
February 750 August 150
March 850 September 250
April 550 October 650
May 350 November 750
June 250 December 830

Manufacturing cost is $200 per set, equally divided between materials and labor. Inventory storage cost is $4 per month. A shortage of sets results in lost sales and is estimated to cost an overall $21 per unit short.

    The inventory on hand at the beginning of the planning period is 200 units. Eight labor hours are required per videodisc player. The workday is seven hours.

Develop an aggregate production schedule for the year using a constant workforce. For simplicity, assume 23 working days each month except July, when the plant closes down for three weeks' vacation (leaving eight working days). Assume that total annual production capacity is greater than or equal to total annual demand (i.e., compute workforce level based on annual demand and annual capacity). (Leave no cells blank - be certain to enter "0" wherever required. Indicate monthly shortages using a negative ending inventory level. Round up the "number of workers" to the next whole number and round down your "monthly production rates" to the next lower whole number.)

January February March April May June July August September October
November
December
Forecast 550 750 850 550 350 250 150 150 250 650 750 830
Beginning inventory
Available production
Ending inventory
Costs Total
Lost sales
Inventory
Total

**Please Help** I having trouble could you show steps please!!
Answer #1

Total annual demand net of beginning inventory = 550+750+850+550+350+250+150+150+250+650+750+830-200 = 5880

Total production hours required = 5880 * 8 hours per unit = 47040 hours

Total number of workdays in the year = 23*11+8 = 261

Annual production hours available per worker = 261*7 = 1827

Constant workforce required = Total production hours required / Production hours availableper worker = 47040/1827 = 26  (rounded-up)

Available Production each month, except July = 26*23*7/8 = 523 (rounded-down)

Available production in July = 26*8*7/8 = 182  (rounded-down)

The aggregate plan based on constant workforce as determined above is following:

N O 3 Beginning inventory 4 Available Production 5 Ending inventory 173 173 851 478 1124 997 770 523 523 523 523 523 5935 523

FORMULAS:

Cell Formula Copy to
C3 =MAX(0,B5) C3:M3
B4 =ROUNDDOWN($B$16*B13/10,0) B4:M4
B5 =B4+B3-B2 B5:M5
B8 =MAX(B5,0)*4 B8:M8
B9 =MAX(-B5,0)*21 B9:M9
B10 =SUM(B8:B9) B10:M10
B13 =B12*7 B13:M13
B15 =(N2-B3)*8
B16 =ROUNDUP(B15/N13,0)
N13 =SUM(B13:M13)

Total cost of the plan = $ 30,468